### v1

@@ 5 Conclusions
We present an approach to combine expert models, based on minimisation of the weighted KL divergence between the experts’ models and a candidate model, that incorporates the agent’s beliefs in the form of expectation constraints. The proposed methodology may be suitable to combine models in a multitude of settings in mathematical finance ranging from climate finance, to volatility modelling, and high-frequency trading. Moreover the resulting optimal model can be reinterpreted as a generative model for, e.g., a dynamic risk-aware reinforcement learning approach to mitigating risks and optimising portfolio allocation [8, 16] or generating climate scenarios that feed into risk management for carbon planning [14].

### v2

@@ 7 Conclusions
We present an approach to combine expert models, based on minimisation of the weighted KL divergence between the experts’ models and a candidate model, that incorporates the agent’s beliefs in the form of expectation constraints. The proposed methodology may be suitable to combine models in a multitude of settings in mathematical finance ranging from climate finance, to volatility modelling, and high-frequency trading. Moreover the resulting optimal model can be deployed as a generative model for, e.g., a dynamic risk-aware reinforcement learning approach to mitigating risks and optimising portfolio allocation [10, 23] or generating climate scenarios that feed into risk management for carbon planning [17].